four particles each of mass m placed at the four corner of square of edge a,move with velocity v,along a circle,which circumscribes the square,under the influence of their mutul gravitational force.find the speed v

Grade:11

3 Answers

Swapnil Saxena
12 years ago

The gravtiatation force on the a single particle is Gm2/a2 + Gm2/a2 + Gm2/2a2 = 5Gm2/2a2

= 5Gm2/2a2= root(2)mv2/a (Centripetal force)

= (5Gm/2a(root(2)))^1/2

 

 


Swapnil Saxena
12 years ago

Sorry for my mistake,

The gravtiatation force on the a single particle is Gm2root(2)/a2 + Gm2/2a= Gm2(2root(2)+1)/2a2

= Gm2(2root(2)+1)/2a2= root(2)mv2/a (Centripetal force)

= (Gm(1+1/2(root(2))/a)^1/2

Aman Bansal
12 years ago

Daer Aashish,

The gravtiatation force on the a single particle is Gm2root(2)/a2 + Gm2/2a= Gm2(2root(2)+1)/2a2

= Gm2(2root(2)+1)/2a2= root(2)mv2/a (Centripetal force)

= (Gm(1+1/2(root(2))/a)^1/2

Best Of luck

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert